# Proactive Recommendation in Social Networks: Steering User Interest via Neighbor Influence

### v1

@@ 4.2.3. Semi-Synthetic Datasets
4.3. Target Users and Target Items Selection
@@
The selection criteria for target user groups are as follows:(1) each user in the target group must have at least one neighbor who is interested in the target item, providing an opportunity for the target user to be steered by their neighbors [^3] PRSN assumes that a user cannot be steered to interact with a target item if none of this user’s neighbors like it.; and (2) individual interest regarding the target item of the user in the group is below a threshold $t\leq 1$, i.e., $\eta_{u^{\ast}i^{\ast}}\leq t$, indicating the user is not very interested in the target item and can to be steered. The selection criteria for a target item is straightforward: once a target item is chosen, the number of qualified target users must exceed a predetermined size. For example, suppose the predetermined size of a target user group is 50, there must be a minimum of 50 users in the entire user set who meet the criteria of target users for the target item.
In practice, we select $1,000$ target items and then select a predetermined size of target user group for each target item [^4] For FilmTrust with items fewer than $1,000$, we choose all possible target items.. We conduct experiments on the aforementioned target user groups and items in turn and average the experiment results to ensure the reliability and effectiveness of our conclusions. The above selection process is implemented by rejecting sampling without replacement. Taking the selection of a target item as an example, we randomly sample an item in the entire item set and determine whether to reject the item according to the criterion for selecting a target item.
4.4. Evaluation Metrics
@@
    (15) $\text{IoIP}\triangleq\frac{1}{|\mathcal{U}^{\ast}|}\sum_{u^{\ast}\in\mathcal{U}^{\ast}}\frac{\sigma(\eta_{u^{\ast}i^{\ast}}+\beta\Delta_{u^{\ast}i^{\ast}})-\sigma(\eta_{u^{\ast}i^{\ast}})}{\sigma(\eta_{u^{\ast}i^{\ast}})}.$
@@
    (16) $\text{DtNE}\triangleq\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}}).$
@@
4.5. Baselines
@@
Existing recommender models are not suitable for PRSN as they do not model the effect of the target item’s exposure to the target user’s neighbors on the target user’s feedback. Therefore, we propose two heuristic strategies to adapt existing recommender models for PRSN as follows:
  • Direct. Expose the target item to the target users’ neighbors who like the item by maximizing the following objective over $\bm{o}_{\mathcal{N}_{u}{i^{\ast}}}$:
    (17) $\begin{split}\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}o_{u^{\prime}i^{\ast}}\left(\sigma\left(\bm{e}^{b}_{u^{\prime}}\cdot\bm{e}^{b}_{i^{\ast}}\right)-0.5\right)-\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}}),\end{split}$
@@
    where $\bm{e}^{b}_{u}$ and $\bm{e}^{b}_{i}$ are the selected recommender model’s user representation and item representation, respectively. This strategy has the same direction of neighbor influence as the simulated user feedback generalization function 13 and 14. For items that are liked, neighbors will have a positive effect on the target user, and vice versa, a negative effect.
  • Sim. Leveraging simulated user feedback generalization function 13 and 14 and maximize it over $\bm{o}_{\mathcal{N}_{u}{i^{\ast}}}$ as follows:
@@
    (18) $\begin{split}\sum_{u^{\ast}\in\mathcal{U^{\ast}}}\sigma\left(\eta\left(\bm{e}^{b}_{u^{\ast}},\bm{e}^{b}_{i^{\ast}}\right)+\beta\Delta\left(\bm{\mathrm{E}}^{b}_{\mathcal{N}_{u^{\ast}}},\bm{e}^{b}_{u^{\ast}},\bm{e}^{b}_{i^{\ast}},\bm{o}_{\mathcal{N}_{u^{\ast}}{i^{\ast}}}\right)\right)-\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c\left(o_{u^{\prime}i^{\ast}}\right),\end{split}$
    where $\bm{\mathrm{E}}^{b}_{\mathcal{N}_{u}}$ is user representation matrix for $\mathcal{N}_{u}$. This strategy has a consistent assumption of neighbor influence with the simulated user feedback generalization function 13 and 14.
@@
Accordingly, we design 6 baselines based on MF, LightGCN(He et al., 2020), and DiffNet (Wu et al., 2019), which are Direct MF, Sim MF, Direct LightGCN, Sim LightGCN, Direct DiffNet, and Sim DiffNet. MF is the most classic and fundamental model in recommendation, which can be viewed as a variant of NIRec removing the aggregation of the item’s exposure to the user’s neighbors and the neighbors’ features. LightGCN is a representative model for graph-based recommendation. DiffNet is the representative model that aggregates the features of the user’s neighbors, which can be viewed as a variant of NeSRec removing the aggregation of the item’s exposure to the user’s neighbors. Note social recommendation methods like DiffNet leverage information from social networks to better align recommendations with users’ interests. However, they fail to steer target users beyond their historical interest because they neglect the effect of exposing an item to neighbors on a user (rely on no-interference assumption).
5. EXPERIMENTS
@@
Table 2. Overall performance without damage constraints across different target user group sizes given $\beta=10$ and $t=1$. The shaded portion is Oracle’s IoIP. The bold and underlined fonts indicate the highest and second-highest IoIP other than Oracle.
Simulation Ciao Sim Epinions Sim Filmtrust Sim
@@
  • RQ4: Can we adjust the target user group to better steer a subset of those users?
To reveal the ideal results, we view the user feedback generation function with simulated representation as a method for PRSN, named Oracle. For brevity, certain experiments’ results about Epinions and Filmtrust are shown in A.
5.1. Overall Performance (RQ1)
5.1.1. Steering without Damage Constraints
@@
Figure 4. DtNE-IoIP curves across different target user group sizes given $\beta=10$ and $t=1$ on Sim Ciao.
5.1.2. Steering with Damage Constraints
@@
  • Without proactive steering (DtNE = 0), IoIP of target user groups with sizes of 50, 75, and 100 (25, 50, and 75 on Sim Filmtrust) are negative. This occurs because the individual interests of the selected target users in the target item are below the threshold ($t\leq 1$), and some users negatively impact others. These results reveal that merely exposing the target item to the target user groups not only fails to improve the interaction probability but may also have a detrimental effect. Therefore, proactive steering is necessary given the existence of neighbor interference.
@@
Figure 5. DtNE-IoIP curves across different interference strength given a group size of 50 and $t=1$ on Sim Ciao.
Figure 6. DtNE-IoIP curves across individual interests threshold given a group size of 50 and $\beta=10$ on Sim Ciao.
Figure 7. NIRec’s DtNE-IoIP curves w.r.t. the original target user group and the adjusted target user groups by different methods across different thresholds given a group size of 50 and $\beta=10$ on Sim Ciao.
5.2. In-Depth Analysis (RQ2, RQ3, RQ4)
5.2.1. Influence of Interference Strength Coefficient (RQ2)
@@
5.2.2. Influence of Individual Interests Threshold (RQ3)
Figure 6 presents DtNE-IoIP curves across different individual interest thresholds given a group size of 50 and $\beta=10$ on Sim Ciao. From Figure 6, we make three key observations: (1) IoIP of all methods decreases as $t$ decreases, showing the difficulty of steering the target user group comprising users who have low interests in the target item; (2) all methods, other than Oracle, exhibit very low IoIP, especially at $t=-1$, where the IoIP is negative across all damage constraints. This reveals that the negative impact produced in the target user group is too large to simultaneously steer users in the group when methods other than Oracle are not ideal; and (3) NIRec outperforms all baselines at $t=0$ and $t=-1$, revealing it is better able to mitigate the negative impact produced in the target group.
5.2.3. Target User Group Adjustment Study (RQ3)
@@
Figure 7 shows NIRec’s DtNE-IoIP curves w.r.t. the original and adjusted target groups using different methods across different thresholds given a group size of 50 and $\beta=10$ in Sim Ciao. In particular, we have two strategies to select users in the target user group: (1) use different methods to respectively select users in the target user group that produce the highest negative impact on interaction probability; and (2) randomly select users in the target user group. The adjusting process is: (1) selecting users in the target user group based on the above strategies; (2) cutting the social edges from the selected users to other users in the group; and (3) comparing IoIP of NIRec on different adjusted target user groups to IoIP of NIRec on the original group. We observe: (1) NIRec’s IoIP significantly improves after adjusting the target user group using any method, indicating that adjustment is important for improving steering performance; (2) adjustments by strategy 1 are better than strategy 2, showing all methods can effectively identify target users’ negative impact in the target group rather than only decreasing social edges; and (3) adjustment by NIRec works best, validating its effectiveness. In real-world applications, cutting social edges is challenging. Nonetheless, these results provide two valuable insights for better steering performance: (1) construct a group where there are fewer social edges between users within the target group, which leads to a lower negative impact produced in the group; and (2) first remove a select number of users producing the highest negative impact, then focus on steering the remaining users in the group, which results in improved performance.
6. RELATED WORKS
6.1. Steering Recommendation
The first line steers users without a target item, which directly improves recommendation diversity (Carbonell and Goldstein, 1998; Chen et al., 2018; Zheng et al., 2021; Wang et al., 2022; Yang et al., 2023b; Zhang et al., 2023b). For example,  (Carbonell and Goldstein, 1998) reduces the redundancy of the recommendation list using greedy search-based re-ranking.  (Zhang et al., 2023b) recommends multiple categories of items to users through disentangled representation learning. While these methods broaden users’ exposure to items beyond their historical interests, they do not account for users’ willingness to engage with those items, which limits their ability to effectively steer user interests. The other line steers with a target item(Zhu et al., 2023; Bi et al., 2024). They sequentially expose a series of intermediate items to a target user, creating a bridge between the user’s existing interests and the target item. However, the alignment of users’ interests with the sequence remains uncertain, especially when there is a substantial divergence between their current interests and the target item. Should users stray from the trajectory, these methods risk failure.
6.2. Causal Recommendation
One widely discussed task is to use a potential outcome framework to address the problem of missing not at random in recommendation (Wang et al., 2019; Dai et al., 2022; Li et al., 2023a, c; Yang et al., 2023a; Li et al., 2023b). These methods rely on the no-interference assumption and focus on the unbiased learning of potential outcomes for all user-item pairs with exposure treatment values of 1.  (Schnabel et al., 2016; Imbens and Rubin, 2015) use various strategies to estimate inverse propensity scores and reweight the biased data to align the distribution of unbiased data.  (Jiang and Li, 2016; Wang et al., 2019; Chen et al., 2021; Dai et al., 2022) and multiple robust (Li et al., 2022) incorporate loss imputation with inverse propensity scores to achieve double or multiple robustness for unbiased recommendation.  (Li et al., 2023c) proposes a stable doubly robust learning method to deal with high variance from small propensity scores.  (Yang et al., 2023a) proposes to learn unbiased representation for the potential outcomes by representation balancing. Causal recommendation also involves other problems with interference assumption (Wu et al., 2022). For example,  (Wu et al., 2022) formulates the effect of position on user feedback within a potential outcome framework by considering the item’s position in the recommendation list as a treatment.  (Sato et al., 2019, 2020; Xie et al., 2021) consider the scenario that a user can interact with an item without recommendation and estimate the uplift of the user’s feedback from recommendation by the potential outcome framework. Unlike them, our work relaxes the no-interference assumption and focuses on modeling interference-aware potential outcomes.
6.3. Causal Inference under Network Interference
Network interference receives widespread attention in causal inference for estimating individual treatment effects (Ogburn and VanderWeele, 2014; Ma and Tresp, 2021; Forastiere et al., 2021; Ma et al., 2022; Jiang and Sun, 2022; Zhao et al., 2022). For example,  (Ogburn and VanderWeele, 2014) distinguishes between different causal mechanisms that give rise to interference using causal graphs.  (Ma et al., 2022) learns interference representation by hypergraph convolution with attention. Considering interference,  (Ma and Tresp, 2021) maximizes the utility on the entire network by optimizing all units’ treatments in the population on the network. Our work differs from these studies in two key ways: (1) constructing target user-item pairs and adjusting the neighbors’ treatment values to achieve our task for those target pairs; and (2) considering user experience under treatment adjusting.
7. CONCLUSION
@@ Appendix A Experiments
In contrast to the results obtained in Sim Ciao, experiments conducted in Sim Epinions and Sim Filmtrust often exhibit positive IoIP. We attribute this phenomenon to the lower social edge density within Epinions and Filmtrust, which consequently leads to lower edge density among users within the target user group. This experimental observation also conveniently corroborates the insight from the target user group adjustment study —constructing a group where there are fewer social edges between users within the group will lead to a lower negative impact produced in the group and achieve better steering performance.
Figure 8. DtNE-IoIP curves across different target user group sizes given $\beta=10$ and $t=1$ in Sim Epinions and Sim Filmtrust.
Figure 9. DtNE-IoIP curves across different interference strength coefficients given a group size of 50 and $t=1$ in Sim Epinions and Sim Filmtrust.
Figure 10. DtNE-IoIP curves across individual interests threshold given a group size of 50 and $\beta=10$ in Sim Epinions and Sim Filmtrust.
Figure 11. NIRec’s DtNE-IoIP curves w.r.t. the original target user group and the adjusted target user groups by different methods across different thresholds given a group size of 50 and $\beta=10$ in Sim Epinions and Sim Filmtrust.

### v2

@@ 5.2.3 Semi-Synthetic Datasets
5.2.4 Discussion for Semi-Simulation
Our semi-simulation aims to address the absence of user feedback under varying neighbor interference in offline datasets. It consists of two components: the mechanism of neighbor influence on user feedback in social networks and the user feedback dataset used for training. For the influence mechanism of neighbors on user feedback, we rely on a well-known and widely accepted assumption: users’ feedback is influenced by their neighbors’ interests and the strength of social relationships [lewis2012social]. Specifically, we assume that recommending the target item to the neighbors of the target user can influence the target user’s feedback on that item, and this influence can be either positive or negative. We implement this mechanism based on real-world offline datasets, where the user/item representation simulation follows the design in [guo2021mrdr], and the user feedback generation function follows [ma2021interference, ma2022hypergraph]. For the user feedback set used for training, we align the semi-synthetic data with real-world offline datasets as much as possible, including the propensity of observing feedback, randomness, data sparsity, and the number of connected users exposed to the same item. These efforts ensure that our semi-simulation reflects real-world scenarios as closely as possible. Nonetheless, our current semi-simulation lacks a mechanism to model how influence propagates through multi-hop neighbors. We plan to address these issues in future work.
5.3 Target Users and Target Items Selection
@@
The selection criteria for target user groups are as follows:(1) each user in the target group must have at least one neighbor who is interested in the target item, providing an opportunity for the target user to be steered by their neighbors [^3] PRSN assumes that a user cannot be steered to interact with a target item if none of this user’s neighbors like it.; and (2) individual interest regarding the target item of the user in the group is below a threshold $t\leq 1$, i.e., $\eta_{u^{\ast}i^{\ast}}\leq t$, indicating the user is not very interested in the target item and can be steered. The selection criterion for a target item is straightforward: once a target item is chosen, the number of qualified target users must exceed a predetermined size. For example, suppose the predetermined size of a target user group is 50, there must be a minimum of 50 users in the entire user set who meet the selection criteria of target users for the target item.
In practice, we select $1,000$ target items and then select a predetermined size of target user group for each target item [^4] For Filmtrust with items fewer than $1,000$, we choose all possible target items.. We conduct experiments on the aforementioned target user groups and items in turn and average the experiment results to ensure the reliability and effectiveness of our conclusions. The above selection process is implemented by rejection sampling without replacement. Taking the selection of a target item as an example, we randomly sample an item in the entire item set and determine whether to reject the item according to the criterion for selecting a target item.
5.4 Evaluation Metrics
@@
    $\text{IoIP}\triangleq\frac{1}{|\mathcal{U}^{\ast}|}\sum_{u^{\ast}\in\mathcal{U}^{\ast}}\frac{\sigma(\eta_{u^{\ast}i^{\ast}}+\beta\Delta_{u^{\ast}i^{\ast}})-\sigma(\eta_{u^{\ast}i^{\ast}})}{\sigma(\eta_{u^{\ast}i^{\ast}})}.$ (15)
@@
    $\text{DtNE}\triangleq\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}}).$ (16)
@@
IoIP directly corresponds to our first goal of PRSN — whether steering raises a target user’s probability of interacting with the target item. Formally, it computes the average relative uplift in interaction probability, normalized by each user’s initial interaction probability without steering. IoIP enjoys the following properties: (1) relativity: by dividing by the initial probability, it fairly compares gains across users with different initial interaction probabilities; (2) boundedness: it has a lower bound (-1) and no upper bound—this means it can strongly reward increasing interaction probabilities for items that initially have low probabilities, precisely matching the proactive recommendation objective of steering users beyond their historical interests; (3) monotonicity: any positive steering adjustment yields a non‑negative increase. DtNE directly corresponds to our second goal—how much steering damages the experience of the neighbors of the target users. It calculates the total damage across all neighbors. DtNE enjoys the following properties: (1) cumulative measurement: it sums the total damage to all neighbors’ experience, enabling holistic auditing; (2) non‑negativity: it is always greater than or equal to 0, so zero steering yields zero damage; (3) flexibility: by selecting different cost functions, one can measure the damage to the neighbors’ experience from different perspectives. The cost function used in this study is defined in Section 4.3.2.
5.5 Baselines
@@
Existing recommender models are not suitable for PRSN as they do not model the effect of the target item’s exposure to the target user’s neighbors on the target user’s feedback. We are the first to do this task. Therefore, we propose two heuristic strategies to adapt existing recommender models for PRSN as follows:
@@
  • Direct. Take the same direction of neighbor interference as the user feedback generation function 13. For items that are liked, neighbors will have a positive effect on the target user, and vice versa. Formally, the strategy is maximizing the following objective over $\bm{o}_{\mathcal{N}_{u}{i^{\ast}}}$:
@@
    $\begin{split}\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}o_{u^{\prime}i^{\ast}}\left(\sigma\left(\bm{e}^{b}_{u^{\prime}}\cdot\bm{e}^{b}_{i^{\ast}}\right)-0.5\right)-\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c(o_{u^{\prime}i^{\ast}}),\end{split}$ (17)
    where $\bm{e}^{b}_{u}$ and $\bm{e}^{b}_{i}$ are the selected recommender model’s user representation and item representation, respectively.
  • Sim. Take a consistent assumption of neighbor interference with user feedback generation function 13. Formally, the strategy is maximizing the following objective over $\bm{o}_{\mathcal{N}_{u}{i^{\ast}}}$:
    $\begin{split}&\sum_{u^{\ast}\in\mathcal{U^{\ast}}}\sigma\left(\eta\left(\bm{e}^{b}_{u^{\ast}},\bm{e}^{b}_{i^{\ast}}\right)+\beta\Delta\left(\bm{\mathrm{E}}^{b}_{\mathcal{N}_{u^{\ast}}},\bm{e}^{b}_{u^{\ast}},\bm{e}^{b}_{i^{\ast}},\bm{o}_{\mathcal{N}_{u^{\ast}}{i^{\ast}}}\right)\right)\\ &-\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c\left(o_{u^{\prime}i^{\ast}}\right),\end{split}$ (18)
    where $\bm{\mathrm{E}}^{b}_{\mathcal{N}_{u}}$ is user representation matrix for $\mathcal{N}_{u}$.
@@
Accordingly, we design 6 baselines based on MF, LightGCN [He2020lightgcn], and DiffNet [wu2019diffnet], which are Direct MF, Sim MF, Direct LightGCN, Sim LightGCN, Direct DiffNet, and Sim DiffNet. MF is the most classic and fundamental model in recommendation, which can be viewed as a variant of NIRec removing the aggregation of the item’s exposure to the user’s neighbors and the neighbors’ features. LightGCN is a representative model for graph-based recommendation. DiffNet is the representative model that aggregates the features of the user’s neighbors, which can be viewed as a variant of NIRec removing the aggregation of the item’s exposure to the user’s neighbors. Note that social recommendation methods like DiffNet leverage information from social networks to better align recommendations with users’ interests. However, they fail to steer target users beyond their historical interest because they neglect the effect of exposing an item to neighbors on a user (rely on no-interference assumption).
6 EXPERIMENTS
@@
Table 2: Overall performance without damage constraints across different target user group sizes given $\beta=10$ and $t=1$. The bold and underlined fonts indicate the highest and second-highest IoIP other than Oracle.
Simulation Sim Ciao Sim Epinions Sim Filmtrust
@@
  • RQ4: Can the effectiveness of steering user interests be further improved by adjusting the target user group?
To reveal the ideal results, we view the user feedback generation function with the simulated representation as a method for PRSN, named Oracle. For brevity, certain experiments’ results about Epinions and Filmtrust are shown in A.
6.1 Overall Performance (RQ1)
6.1.1 Steering without Damage Constraints
@@
Figure 4: DtNE-IoIP curves across different target user group sizes given $\beta=10$ and $t=1$.
6.1.2 Steering with Damage Constraints
@@
  • Without proactive steering (DtNE = 0), IoIP of target user groups with sizes of 50, 75, and 100 (25, 50, and 75 on Sim Filmtrust) are negative. This occurs because the individual interests of the selected target users in the target item are below the threshold ($t\leq 1$), leading to scenarios where these users negatively influence each other’s interest in the target item through the social network. These results reveal that merely exposing the target item to the target user groups not only fails to improve the interaction probability but may also have a detrimental effect. Therefore, proactive steering is necessary, given the existence of neighbor interference.
@@
Figure 5: DtNE-IoIP curves across different interference strength given a group size of 50 and $t=1$ in Sim Ciao.
Figure 6: DtNE-IoIP curves across individual interests threshold given a group size of 50 and $\beta=10$ in Sim Ciao.
Figure 7: NIRec’s DtNE-IoIP curves w.r.t. the original target user group and the adjusted target user groups by different methods across different thresholds given a group size of 50 and $\beta=10$ in Sim Ciao.
6.2 In-Depth Analysis (RQ2, RQ3, RQ4)
6.2.1 Influence of Interference Strength Coefficient (RQ2)
@@
6.2.2 Influence of Individual Interests Threshold (RQ3)
Figure 6 presents DtNE-IoIP curves across different individual interest thresholds given a group size of 50 and $\beta=10$ on Sim Ciao. In Figure 6, we make three key observations:
  • The IoIP of all methods decreases as $t$ becomes smaller, indicating that the lower the target group’s interest in the target item, the harder it is to steer users within the target group.
  • All methods, other than Oracle, exhibit very low IoIP, especially at t=1t=-1, where the IoIP is negative across all damage constraints. This indicates that when the interference strength estimation is insufficiently accurate, we cannot simultaneously steer all users within the target user group who have extremely low interest in the target item. The reason is that when target users have very low interest in the target item, the negative effects associated with the target item, propagated through the social network among users within the group, may outweigh the positive effects of steering efforts.
  • NIRec outperforms all baselines at $t=0$ and t=1t=-1, demonstrating its ability to better mitigate the negative impact on the target group.
6.2.3 Target User Group Adjustment Study (RQ4)
@@
Figure 7 shows NIRec’s DtNE-IoIP curves w.r.t. the original and adjusted target groups using different methods across different thresholds given a group size of 50 and $\beta=10$ in Sim Ciao. In particular, we have two strategies to select users in the target user group: (1) use different methods to respectively select users in the target user group that produce the highest negative impact on interaction probability; and (2) randomly select users in the target user group. The adjusting process is: (1) selecting users in the target user group based on the above strategies; (2) cutting the social edges from the selected users to other users in the group; and (3) comparing IoIP of NIRec on different adjusted target user groups to IoIP of NIRec on the original group. We observe: (1) NIRec’s IoIP significantly improves after adjusting the target user group using any strategy and any method, indicating that the adjustment of cutting edges in the target user group is important for improving steering performance; (2) adjustments by strategy 1 are better than strategy 2, showing all methods can effectively identify target users’ negative impact in the target group rather than only decreasing the number of social edges; and (3) adjustment by NIRec works best, validating its effectiveness. In real-world applications, cutting social edges is challenging. Nonetheless, these results provide two valuable insights for better steering performance: (1) construct a group where there are fewer social edges between users within the target user group, which leads to a lower negative impact produced in the group; and (2) first remove a select number of users producing the highest negative impact, then focus on steering the remaining users in the target user group, which results in improved performance.
7 CONCLUSION
@@ Appendix A Experiments
In contrast to the results obtained on Sim Ciao, experiments conducted in Sim Epinions and Sim Filmtrust (shown in Figure 8, Figure 9, and Figure 10) often exhibit positive IoIP. We attribute this phenomenon to the lower social edge density within Epinions and Filmtrust, which consequently leads to lower edge density among users within the target user group and a weaker negative influence propagated among users within the target group through the social network. This experimental observation also conveniently corroborates the insight from the target user group adjustment study — constructing a group where there are fewer social edges between users within the group will lead to a lower negative impact produced in the group and achieve better steering performance.
Figure 8: DtNE-IoIP curves across different interference strength coefficients given a group size of 50 and $t=1$ in Sim Epinions and Sim Filmtrust.
Figure 9: DtNE-IoIP curves across individual interests threshold given a group size of 50 and $\beta=10$ in Sim Epinions and Sim Filmtrust.
Figure 10: NIRec’s DtNE-IoIP curves w.r.t. the original target user group and the adjusted target user groups by different methods across different thresholds given a group size of 50 and $\beta=10$ in Sim Epinions and Sim Filmtrust.